OpenStudy (anonymous):

Which of the following is the solution to the equation 8(y – 3) = 2y ?

OpenStudy (anonymous):

y = 1.5 y = 4.5 y = –4.5 y = –1.5

OpenStudy (hoblos):

8(y – 3) = 2y 8y -24 = 2y 6y = 24 y=4

OpenStudy (hoblos):

not one of the options though !!!

OpenStudy (anonymous):

yeahhh. i dont think 4 is the answer :(

OpenStudy (hoblos):

is it just like this "8(y – 3) = 2y" or there is something missing ??

OpenStudy (anonymous):

okay, i think i wrote it wrong. it should be 8^(y-3) =2^y

OpenStudy (anonymous):

Use logaritms in order to solve that.

OpenStudy (hoblos):

0k this makes a lot of difference...... 8^(y-3) =2^y remember that 8=2^3 .. 2^3(y-3) = 2^y then 3(y-3) = y 3y-9 = y 2y = 9 y=4.5

OpenStudy (anonymous):

thank you so so much.!! :)

OpenStudy (anonymous):

Just think it through, you can't use 1.5 or -1.5, or else it will make the equation false. You are left with 4.5 and -4.5. -4.5 won't work since you will end up with a large negative number compared to a negative number.

OpenStudy (anonymous):

(y-3)*ln(8)=ln(2)*y y*ln(8)-3*ln(8)=ln(2)*y y(ln(8)-ln(2))=3ln(8) y=3ln(8)/(ln(8)-ln(2)) ---> y=9/2=4.5